Find the orthogonal trajectory curve of all the lines passing through the origin that is y = C


1
Expert's answer
2021-11-15T16:24:30-0500

Let the equation be "y = Cx," where "C" is an arbitrary constant.

Differentiating with respect to "x," we get, "y'=C."

Eliminate "C"


"y'=\dfrac{y}{x}"

which is the differential equation of a family of lines. 

Replace "y'" with "(-1\/y')"


"-\dfrac{1}{y'}=\dfrac{y}{x}"

"ydy=-xdx"

Integrate


"\int ydy=-\int xdx"

"\dfrac{y^2}{2}=-\dfrac{x^2}{2}+c_1"

"x^2+y^2=2c_1"

By replacing "2c_1" with "R^2" we see that the orthogonal trajectories for the family of straight lines are concentric circles 

"x^2+y^2=R^2"